OpenStudy (anonymous):

Solve logx + 5 36 = 2

OpenStudy (campbell_st):

is the equation \[\log x + 536 = 2\]

OpenStudy (anonymous):

my guess is it is \[\large \log_{x+5}(36)=2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

write in equivalent exponential form as \((x+5)^2=36\) and solve that one

OpenStudy (campbell_st):

raise each side as powers of (x +5) then 36 = (x + 5) ^2 and then solve the quadratic

OpenStudy (anonymous):

takes two steps \[(x+5)^2=36\] \[x+5=6\] \[x=1\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

actually you should get two solutions to \[(x+5)^2=36\] but you can ignore the negative solution because the base of a logarithmic function is never negative